Dario puts 44 marbles in a box in which 14 are red, 12 are blue, 18 are yellow. If Dario picks one marble at random, what is the probability that he selects a red marble or a yellow marble ?
Found 2 solutions by jim_thompson5910, Edwin McCravy:
Answer by jim_thompson5910(35256)

There are 14+18 = 32 marbles that are either red or yellow.

This is out of 44 marbles total.

Divide the two values (32 and 44) to get: 32/44

Now reduce: 32/44 = 8/11

To reduce, I divided both parts by the GCF 4

The final answer as a fraction is 8/11

Use a calculator to find that 8/11 = 0.7273 approximately

The approximate answer in decimal form is 0.7273 (which represents roughly 72.73% probability)

Answer by Edwin McCravy(20055)